# From Neck to Head: Bio-Impedance Sensing for Head Pose Estimation

### v1

@@ Abstract.
We present NeckSense, a novel wearable system for head pose tracking that leverages multi-channel bio-impedance sensing with soft, dry electrodes embedded in a lightweight, necklace-style form factor. NeckSense captures dynamic changes in tissue impedance around the neck, which are modulated by head rotations and subtle muscle activations. To robustly estimate head pose, we propose a deep learning framework that integrates anatomical priors, including joint constraints and natural head rotation ranges, into the loss function design. We validate NeckSense on 7 participants using the current SOTA pose estimation model as ground truth. Our system achieves a mean per vertex error of 25.9 mm across various head movements with a leave one person out cross validation method, demonstrating that a compact, line-of-sight–free bio-impedance wearable can deliver head tracking performance comparable to state-of-the-art vision-based methods.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06submissionid: 8747ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
We proposed the NeckSense: a lightweight necklace integrating soft, dry, and reusable electrodes for comfortable, continuous wear. The hardware is paired with a transformer-based deep learning framework, Imp2Head, which maps the raw impedance signals to 3D head pose. Critically, our model incorporates anatomical constraints, such as natural joint limits, to ensure its predictions are physically plausible and robust. To validate NeckSense’s performance in realistic conditions, we conducted a user study (N=7). The protocol was specifically designed to capture a diverse and representative range of motions found in naturalistic scenarios like conversation and AR/VR interaction, rather than limiting participants to scripted gestures. We used a state-of-the-art vision-based model for ground-truth labeling. Our system achieved a mean per-vertex error of 25.9 mm, a result comparable to the vision-based system’s own inherent error, demonstrating that the Necksense is a competitive alternative.
@@ 2. Related Work
Wearable pose estimation typically employs compact sensor modalities, including IMU (Kim et al., 2024; Georgi et al., 2015; Tavakoli et al., 2018; Mollyn et al., 2023; Xu et al., 2024), capacitive sensors (Ziraknejad et al., 2014; Bian et al., 2024; Zhou et al., 2023), bio-transducers, like EMG (Georgi et al., 2015; Aguiar and Bó, 2017; Tavakoli et al., 2018; Meta Reality Labs, 2025; Chen et al., 2021) and acoustical sensors (Iravantchi et al., 2019), due to their small form factor and suitability for body-mounted applications. IMUs estimate head orientation by integrating angular velocity and acceleration signals (Leonardis et al., 2022). Single-IMU setups, typically integrated into headbands or eyeglasses, are simple but susceptible to drift (Severin, 2020). Multi-IMU arrays improve accuracy but add complexity, bulk, and discomfort (Breen et al., 2009). Hybrid IMU-camera approaches partially resolve drift but are heavily influenced by environmental factors such as lighting and visibility. EMG-based systems effectively track gestures and limb movements by capturing muscle activation signals (Georgi et al., 2015; Aguiar and Bó, 2017). However, EMG inherently relies on active muscle contraction, limiting its efficacy for continuous or static head pose estimation. Recent EMG-based head tracking systems necessitate electrodes on neck and facial areas, significantly reducing comfort and social acceptability (Kim et al., 2024). Additionally, variations in electrode placement can severely affect long-term reliability. Capacitive sensing offers another promising alternative for gesture and pose estimation (Bian et al., 2024, 2022), yet its susceptibility to environmental interference limits its robustness in everyday use cases (Bian et al., 2024). Thus, existing wearable sensor solutions face inherent trade-offs among accuracy, portability, usability, and environmental robustness, motivating continued exploration of complementary sensing modalities.
@@
Bio-impedance sensing has been extensively explored for various physiological monitoring tasks, such as respiration (Jeong et al., 2019), hydration (Sato et al., 2017), and swallowing detection (Lin et al., 2019). More recently, bio-impedance has emerged as a promising modality for gesture recognition in human-computer interaction (Waghmare et al., 2023; Zhang and Harrison, 2015; Liu et al., 2025) and activity recognition (Liu et al., 2024b, c, a). Unlike EMG, bio-impedance does not require active muscle contraction, enabling continuous detection of both static and dynamic postures based on tissue geometry changes alone (Sanchez and Rutkove, 2017; Stupin et al., 2021). Furthermore, bio-impedance inherently supports multi-channel volumetric sensing, capturing spatial deformation patterns across body regions to provide richer contextual information than localized EMG or angular IMUs (Zhang and Harrison, 2015).
@@
EIT, a specific bio-impedance approach, typically reconstructs internal conductivity distributions using boundary voltage and current measurements, requiring numerous electrodes (often more than eight) to achieve high-resolution imaging (Zhang and Harrison, 2015; Zhang et al., 2016; Kyu et al., 2024). Traditional EIT employs a pairwise adjacent current stimulation pattern (Sheffield method), resulting in increased hardware complexity and data processing burdens due to the large number of measurements (specifically, $N(N-3)/2$ for an $N$-electrode system). Additionally, EIT’s inverse reconstruction is highly sensitive to noise, often leading to significant artifacts.
In contrast to traditional EIT, our approach, NeckSense, argues against the necessity of full tomographic reconstruction for effective head pose estimation. Our method employs a significantly simpler hardware setup, using one electrode for stimulation and four additional electrodes placed strategically on the neck to measure the response. Instead of reconstructing tomographic images, our system directly leverages frequency-domain features, signal magnitude and phase, to estimate head pose efficiently and robustly. This approach simplifies the system architecture, reduces computational complexity, and enhances resistance to measurement noise, thus distinguishing itself clearly from conventional EIT methods.
@@
By situating NeckSense within existing bio-transducer literature and explicitly clarifying its unique differences from traditional EIT, we highlight the physiological distinctiveness, practical advantages, and novelty of our simplified bio-impedance sensing approach for wearable head pose estimation and human-computer interaction.
@@ 3.1. General Principle
Figure 1. General principle of bio-impedance sensing-based head pose estimation (Five electrodes integrated in the necklace are attached to the neck to monitor the neck impedance variation caused by head movement, the bio-impedance depends on tissue composition, geometry and distance between the electrodes as well as the orientation of muscle fibers or structure deformation)
@@ 3.2. Hardware Implementation
Figure 2. NeckSense system Design. The system consists of five soft, dry, and reusable electrodes positioned around the neck, with one electrode serving as a common electrode to generate the stimuli. Impedance is measured across four channels, each capturing both magnitude and phase information. The electrodes require no adhesive, offering improved comfort and reusability. Impedance signals are acquired using the AD5941 analog front-end, controlled by an ESP32-S2 microcontroller. Measurement data is wirelessly transmitted to a computer via Bluetooth for real-time processing.
@@ 4.3. Evaluation Metric
[…]
@@ 5. Discussion and Limitation
Our evaluation demonstrates the potential of bio-impedance sensing as a robust, non-intrusive method for wearable head pose estimation. Imp2Head significantly reduces the MPJPE and MPVE compared to baseline models, highlighting the advantage of integrating anatomical biomechanics constraints. Unlike traditional vision-based methods, our impedance-based approach is unaffected by occlusions or varying lighting conditions, making it highly suitable for diverse environments and daily use. The corrected MPJPE of 25.9 mm, while slightly higher than the OS-X vision-based model (24.9 mm), showcases the practical viability of bio-impedance as a competitive alternative. In addition, bio-impedance sensing shows promising potential for jaw pose estimation, a challenging task for other modalities (e.g., IMU, EMG, and capacitive sensors), typically limited to neck-mounted placement. As demonstrated in Fig. 4, our neck-impedance-based model accurately tracks both head and jaw poses, closely following the ground truth trajectories. This highlights bio-impedance’s unique capability to capture subtle anatomical movements beyond the scope of conventional wearable sensors.
Despite the promising results, several limitations of this work warrant discussion. First, the system is trained and evaluated using a vision-based pose estimator (OS-X) rather than a gold-standard motion capture (MoCap) system. While convenient, this reliance on pseudo-ground truth introduces potential bias and limits the upper-bound accuracy of our model. Any inaccuracies or drift in the vision pipeline may propagate through training and evaluation, potentially masking failure cases that would emerge with MoCap-based ground truth. Second, although the use of dry, soft electrodes enhances comfort and wearability, the system remains sensitive to minor shifts in electrode placement, caused by neck movement, sweat, or collar adjustment, which can introduce signal noise and drift. This spatial sensitivity may affect robustness in real-world deployments, especially under long-term use or inconsistent device reattachment, underscoring the need for adaptive preprocessing, auto-calibration, or self-alignment mechanisms. Third, the model does not incorporate individual anatomical differences (e.g., neck size, posture habits, or tissue composition), which could affect impedance patterns. Integrating personalized calibration or subject-specific modeling may further enhance performance across diverse users. Furthermore, the limited sample size (N=7) presents a constraint, potentially affecting the robustness and statistical power of the findings.
@@ 6. Conclusion
In this work, we introduced NeckSense, a novel bio-impedance-based wearable system designed for continuous and unobtrusive head pose estimation. NeckSense leverages multi-channel impedance measurements with soft, dry and reusable electrodes integrated into a lightweight, necklace-style device. Our Transformer-based Imp2Head model effectively captures the relationship between impedance signals and head pose, enhanced by incorporating biomechanical constraints. Experimental evaluation demonstrates that our approach achieves accuracy comparable to state-of-the-art vision-based systems, highlighting the effectiveness and practicality of bio-impedance sensing as an alternative to traditional methods. Future research directions include addressing individual anatomical variability and extending the method to more dynamic motion scenarios. Overall, NeckSense represents a significant advancement in wearable sensing technology, opening new avenues for unobtrusive, reliable, and versatile human pose tracking applications.

### v2

@@ Abstract.
We present NeckSense, a novel wearable system for head pose tracking that leverages multi-channel bio-impedance sensing with soft, dry electrodes embedded in a lightweight, necklace-style form factor. NeckSense captures dynamic changes in tissue impedance around the neck, which are modulated by head rotations and subtle muscle activations. To robustly estimate head pose, we propose a deep learning framework that integrates anatomical priors, including joint constraints and natural head rotation ranges, into the loss function design. We validated NeckSense on 7 participants using the current state-of-the-art pose estimation model as ground truth. Our system achieves a mean per-vertex error of 25.9 mm across various head movements with a leave-one-person-out cross-validation method, demonstrating that a compact bio-impedance wearable can deliver head-tracking performance comparable to state-of-the-art vision-based methods.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 2025 ACM International Symposium on Wearable Computers; October 12–16, 2025; Espoo, Finlandbooktitle: Proceedings of the 2025 ACM International Symposium on Wearable Computers (ISWC ’25), October 12–16, 2025, Espoo, Finlanddoi: 10.1145/3715071.3750425isbn: 979-8-4007-1481-8/2025/10ccs: Human-centered computing Ubiquitous and mobile computing systems and tools
@@ 1. Introduction
We present NeckSense, a lightweight necklace with soft, dry, reusable electrodes for comfortable, continuous wear. Paired with a transformer-based model, Imp2Head, it maps raw impedance signals to 3D head pose while incorporating anatomical constraints to ensure physically plausible predictions. To evaluate performance, we conducted a user study (N=7) using naturalistic motion scenarios like conversation and AR/VR interaction, with ground truth from a state-of-the-art vision-based model. NeckSense achieved a mean per-vertex error of 25.9 mm, comparable to the vision system’s own error, demonstrating it as a competitive alternative.
@@ 2. Related Work
Wearable pose estimation typically employs compact sensor modalities, including IMU (Kim et al., 2024; Georgi et al., 2015; Tavakoli et al., 2018; Mollyn et al., 2023; Xu et al., 2024), capacitive sensors (Ziraknejad et al., 2014; Bian et al., 2024; Zhou et al., 2023), bio-transducers, like EMG (Georgi et al., 2015; Aguiar and Bó, 2017; Tavakoli et al., 2018; Meta Reality Labs, 2025; Chen et al., 2021) and acoustical sensors (Iravantchi et al., 2019), due to their small form factor and suitability for body-mounted applications. IMUs estimate head orientation by integrating angular velocity and acceleration signals (Lopez-Nava and Munoz-Melendez, 2016). Single-IMU setups, typically integrated into headbands or eyeglasses, are simple but susceptible to drift (Severin, 2020). Multi-IMU arrays improve accuracy but add complexity, bulk, and discomfort (Breen et al., 2009). Hybrid IMU-camera approaches partially resolve drift but are heavily influenced by environmental factors such as lighting and visibility. EMG-based systems effectively track gestures and limb movements by capturing muscle activation signals (Georgi et al., 2015; Aguiar and Bó, 2017). However, EMG inherently relies on active muscle contraction, limiting its efficacy for continuous or static head pose estimation. Recent EMG-based head tracking systems necessitate electrodes on neck and facial areas, significantly reducing comfort and social acceptability (Kim et al., 2024). Additionally, variations in electrode placement can severely affect long-term reliability. Capacitive sensing offers another promising alternative for gesture and pose estimation (Bian et al., 2024, 2022), yet its susceptibility to environmental interference limits its robustness in everyday use cases (Bian et al., 2024). Thus, existing wearable sensor solutions face inherent trade-offs among accuracy, portability, usability, and environmental robustness, motivating continued exploration of complementary sensing modalities.
@@
Bio-impedance sensing has been extensively explored for various physiological monitoring tasks, such as respiration (Liu et al., 2025), hydration (O’brien et al., 2002), and swallowing detection (Kusuhara et al., 2004). More recently, bio-impedance has emerged as a promising modality for gesture recognition in human-computer interaction (Waghmare et al., 2023; Zhang and Harrison, 2015; Liu et al., 2025) and activity recognition (Liu et al., 2024b, c, a). Unlike EMG, bio-impedance does not require active muscle contraction, enabling continuous detection of both static and dynamic postures based on tissue geometry changes alone (Sanchez and Rutkove, 2017; Stupin et al., 2021). Furthermore, bio-impedance inherently supports multi-channel volumetric sensing, capturing spatial deformation patterns across body regions to provide richer contextual information than localized EMG or angular IMUs (Zhang and Harrison, 2015).
@@
EIT, a specific bio-impedance approach, typically reconstructs internal conductivity distributions using boundary voltage and current measurements, requiring numerous electrodes (often more than eight) to achieve high-resolution imaging (Zhang and Harrison, 2015; Zhang et al., 2016; Kyu et al., 2024). Traditional EIT employs a pairwise adjacent current stimulation pattern (Sheffield method), resulting in increased hardware complexity and data processing burdens due to the large number of measurements (specifically, N(N3)/2N(N-3)/2 for an $N$-electrode system). Additionally, EIT’s inverse reconstruction is highly sensitive to noise, often leading to significant artifacts.
@@
Unlike traditional EIT, NeckSense eliminates the need for full tomographic reconstruction in head pose estimation. It uses a simplified setup—one stimulation electrode and four strategically placed measurement electrodes, while directly leveraging frequency domain features (magnitude and phase) for efficient, robust pose estimation. This reduces hardware complexity, computational load, and sensitivity to noise. By situating NeckSense within existing bio-transducer literature and explicitly clarifying its unique differences from traditional EIT, we highlight the physiological distinctiveness, practical advantages, and novelty of our simplified bio-impedance sensing approach for wearable head pose estimation and human-computer interaction.
@@ 3.2. Hardware Implementation
Figure 1. Bio-impedance-based head pose estimation uses five necklace-integrated electrodes to monitor neck impedance changes caused by head movement, which depend on tissue composition, electrode geometry and spacing, and muscle orientation or structural deformation
Figure 2. NeckSense uses five soft, reusable electrodes around the neck, with one for stimulation and four for impedance sensing (magnitude and phase). Signals from the AD5941 are read by an ESP32-S2 and sent via Bluetooth for real-time processing
@@ 4.3. Evaluation Metric
where $\mathbf{p}_{b,t,j}\in\mathbb{R}^{3}$ is the 3D position of joint $j\in\{\text{neck, head, jaw}\}$.
@@ 5. Discussion and Limitation
Our evaluation highlights bio-impedance sensing as a robust, non-intrusive method for wearable head pose estimation. Imp2Head significantly reduces MPJPE and MPVE compared to baselines, benefiting from anatomical constraint integration. Unlike vision-based methods, our approach is resilient to occlusion and lighting changes, making it well-suited for daily use. With a corrected MPJPE of 25.9 mm, comparable to OS-X’s 24.9 mm, bio-impedance proves to be a viable alternative. Moreover, it enables jaw pose estimation, a task difficult for IMU, EMG, or capacitive sensors. As shown in Fig. 4, our model tracks both head and jaw motion accurately, demonstrating bio-impedance’s unique ability to capture fine anatomical movements beyond the reach of conventional wearables.
Despite promising results, several limitations remain. First, the system relies on a vision-based pose estimator (OS-X) for training and evaluation instead of a gold-standard MoCap system. This introduces potential bias, as errors or drift in the vision pipeline may propagate through training and obscure failure cases. Second, while soft, dry electrodes improve comfort, the system is sensitive to small shifts in placement from neck movement, sweat, or collar adjustments, leading to signal noise and drift. This highlights the need for adaptive preprocessing, auto-calibration, or self-alignment. Third, the model does not account for individual anatomical differences (e.g., neck size, posture, tissue composition), which may affect impedance patterns. Incorporating personalized calibration could improve generalization. Lastly, the small sample size (N=7) limits the robustness and statistical power of the results.
@@ 6. Conclusion
We introduce NeckSense, a bio-impedance-based wearable system for continuous, unobtrusive head pose estimation. It uses multi-channel impedance measurements via soft, dry, reusable electrodes embedded in a lightweight necklace. Our Transformer-based Imp2Head model maps signals to head pose, incorporating biomechanical constraints for enhanced accuracy. Experiments show performance comparable to state-of-the-art vision-based systems, validating bio-impedance as a practical alternative. Future work includes adapting to individual anatomical differences and handling more dynamic motions. Overall, NeckSense advances wearable sensing, enabling reliable and versatile human pose tracking.
Acknowledgements.
This work is funded by the European Union under the EIC Pathfinder programme, with the project 101162257 STELEC.